# A mechanism-informed reinforcement learning framework for shape optimization of airfoils

### v1

@@ 5.3 Reward function
In the context of shape optimal design, the reward function must match the optimization objective. For instance, consider the objective of generating a shape with the minimum drag, denoted as $D(\{sp_{i}\}_{i=0}^{L+1})$, where $\{sp_{i}\}_{i=0}^{L+1}$ represents the sample points of the airfoil with minimal drag. This relationship is described by the following equation:
$\begin{array}[]{cc}&D(\{sp^{*}_{i}\}_{i=0}^{L+1})\leq D(\{sp_{i}\}_{i=0}^{L+1})\qquad\forall sp_{i}\in\mathbb{R}^{2}\\ &\{sp^{*}_{i}\}_{i=0}^{L+1}\equiv\{sp_{i}\}_{i=0}^{L+1}\quad\text{if\leavevmode\nobreak\ }D(\{sp_{i}\}_{i=0}^{L+1})\leq D(\{sp^{*}_{i}\}_{i=0}^{L+1})\end{array}$ (27)
To simplify, we denote the drag at time step $t$ as $D_{t}$. The reward function at time step $t$ is then defined as:
$R_{t}=D_{t-1}-D_{t}.$ (28)
Intuitively, the reinforcement learning process aims to maximize the cumulative reward $\sum_{t=1}^{T}R_{t}$. A pertinent question arises as to whether this optimization goal is equivalent to the original objective. If we define $\{sp^{**}_{i}\}_{i=0}^{L+1}$ as the sample points of the airfoil yielding the maximal reward, then:
$D_{0}+(-D(\{sp^{**}_{i}\}_{i=0}^{L+1}))\geq D_{0}+(-D(\{sp_{i}\}_{i=0}^{L+1}))\qquad\forall sp_{i}\in\mathbb{R}^{2}.$ (29)
@@
By definition, $\{sp^{**}_{i}\}_{i=0}^{L+1}$ is equivalent to $\{sp^{*}_{i}\}_{i=0}^{L+1}$, demonstrating the equivalence of the reinforcement optimization objective to the primary task. A similar proof, such as maximizing the lift-drag ratio, can also be derived in this manner.
In our shape optimization tasks, the optimization process incorporates additional components beyond the straightforward maximization of cumulative rewards. As mentioned in the previous section, the discount factor $\gamma$ plays a pivotal role in calculating the cumulative reward. The discount factor ensures that the optimization process remains focused on achievable improvements, preventing the pursuit of impractically distant objectives.
@@
To further refine the optimization strategy, regularization terms can be added to the reward function. These terms are designed to encourage exploration, a critical aspect of overcoming local optima and discovering globally optimal solutions. By incorporating exploration-driven regularization, the reward function can be represented as:
$R_{t}=(D_{t-1}-D_{t})+\lambda^{\epsilon}_{E}\cdot(D_{0}-D_{t}),$ (32)
@@
where $\lambda^{\epsilon}_{E}$ is a weighting coefficient that balances the original objective with the exploration incentive. When the current shape yields an improvement in the objective function value, the agent gets reward. Conversely, a penalty is incurred for generating a geometry that does not meet the expected criteria. The parameter $\lambda^{\epsilon}_{E}$ shall be decreased with the update of noise. Thus the reward function will encourage exploration for the initial period while it converges to the equation (28) with the iteration get processed. Theoretically, the result shall be equivalent to the primal objective. These mechanisms ensure a more robust and effective search for the optimal airfoil shape, facilitating the discovery of innovative designs that minimize drag while navigating the complex landscape of possible configurations.
5.4 Twin delayed DDPG
@@
With the introduced modules above, we are going to develop our framework to a more robust version, Twin Delayed Deep Deterministic Policy Gradient, which is an algorithm that builds upon the Deep Deterministic Policy Gradient framework. It introduces several key improvements that enhance stability and performance in reinforcement learning tasks. One of the primary advantages of Twin delayed DDPG is its twin-critic network architecture. This design involves using two separate critic networks to estimate the Q-values, and the smaller of the two Q-value estimates is used for the policy update. This approach effectively reduces the overestimation bias often seen in DDPG, leading to more stable and reliable learning.
Additionally, Twin delayed DDPG introduces policy update delays, where the policy network is updated less frequently than the critic networks. This delay helps mitigate the risk of the policy network exploiting inaccuracies in the Q-value estimates, further enhancing the stability of the learning process. Twin delayed DDPG also applies target policy smoothing, adding noise to the target action, which smoothens the value estimate and prevents the policy from exploiting Q-function errors in sharp transitions.
With this more stable reinforcement learning framework, we constructed the algorithm as Algorithm 3.
@@
Data: Initial wing shape data, actor parameters $\theta^{a}$, critic parameters $\theta^{c_{1}}$ and Twin critic parameters $\theta^{c_{2}}$, target networks $\theta^{a}_{target}$, $\theta_{target}^{c_{1}}$ and $\theta_{target}^{c_{1}}$, ReplayBuffer
Result: Trained Twin Delayed DDPG agent
1 Set the target parameters $\theta^{a}_{target}\leftarrow\theta^{a},\theta^{c_{1}}_{target}\leftarrow\theta^{c_{1}},\theta^{c_{2}}_{target}\leftarrow\theta^{c_{2}}$ for each episode in max_episodes do
2       if Exploration or reach the minimum threshold then
3            Reset the environment and get the initial state
4       end if
5      for step in episode do
6             if Exploration then
7                   The agent produces a random action for execution
8             end if
9            else
10                  The agent selects an action based on $\theta^{a}_{target}$ Noises are added to the action and corresponding parameters The agent executes the action based on the current state Update the cache state
11             end if
12            Observe the new state and calculate the reward based on AFVM4CFD;
13             if reward $<min\_reward$ then
14                   Reload the cache state Execute the opposite action with corresponding parameters
15             end if
16            Save the tuple (state, action, reward, next_state) to the ReplayBuffer;
17             if ReplayBuffer has enough experiences then
18                   Sample experiences (Recent, Best, Random) from ReplayBuffer;
19                   Compute the target actions based on $\theta^{a}_{target}$;
20                   Compute the target Q-value $\min(Q_{\theta_{target}^{c_{1}}},Q_{\theta_{target}^{c_{2}}})$;
21                   Update the critic networks $\theta^{c_{1}}$, $\theta^{c_{2}}$ ;
22                   Soft update critic networks $\theta_{target}^{c_{i}}\leftarrow\theta_{target}^{c_{i}}*\tau+(1-\tau)*\theta^{c_{i}}$ for $i=1,2$ if $step\equiv 0\pmod{delay\_step}$  then
23                         Update the actor network $\theta^{a}$ ;
24                         Soft update actor networks $\theta_{target}^{a}\leftarrow\theta_{target}^{a}*\tau+(1-\tau)*\theta^{a}$
25                   end if
26
27             end if
28            Update the state;
29
30       end for
31
32 end for
@@ 6 Numerical Result
6.1 Minimize drag optimization
The initial example delves into minimizing drag calculations, a process whose precision is bolstered by the DWR-based mesh adaptation method. The study utilizes a NACA0012 airfoil, positioned within a circular domain with a radius of $35$, under a boundary condition of mirror reflection. The operational conditions are set with a Mach number of $0.85$ and an attack angle of $0^{\circ}$.
Figure 10: Left: the record curve of noise coefficient. Right: the record curve of drag during the optimization process.
To ensure the integrity of the airfoil’s structure during training, a constraint is imposed, mandating a minimum thickness of 0.1 for sections where $x<0.1$ and $x>0.7$. Should a potential action result in an airfoil design adhering to this thickness constraint, it is deemed physically infeasible, consequently not executed, and results in the issuance of a negative reward.
As delineated in Algorithm 3, the agent embarks on accumulating experience through random actions performed four times, each starting from the baseline configuration and geometry across $64$ time steps. The initial recorded drag value stands at $0.0452106$.
@@ 6.1 Minimize drag optimization
Figure 11: Left: the airfoil generated from the 15th episode. Right: the result of the optimized airfoil.
Training commences with a moderate batch size of $256$, spanning a total of $75$ epochs, with each epoch comprising 8 time steps. The noise coefficient undergoes adjustment every three epochs, following an exponential decline. As illustrated in Figure 10, the drag value exhibits a decrement trend as iterations advance. Even though the record curve is not smooth while the drag value gets increased at some moment, it still evolved towards the optimization objective. The non-smooth phenomenon originates from the fact of noise. Even if the potential action from the actor neural network generates an effective action with corresponding parameters, the noise may distract the agent from the potential actions. However, such a noise module is indispensable to help the agent explore the possibility of different states.
Throughout the optimization process, the agent probes various airfoil shapes. For instance, Figure 11 showcases a configuration from the $15^{th}$ epoch, sporting a drag value of $0.031537$. The drag value’s trajectory halts around $0.02$, constrained by the minimum thickness stipulation. Encounters with this boundary condition trigger a negative reward, steering exploration toward alternate states. Notably, as the exploration progresses, particularly in the concluding epoch, the airfoil’s tail approaches the thickness restriction, a direct consequence of rewards associated with thickness reduction.
The reduction in drag can be elucidated by examining the Mach isolines. Initially, the NACA 0012 airfoil exhibited strong shock waves on both sides of the wing, typically associated with significant pressure jumps near these shocks, contributing to a large value of drag. However, the optimized airfoil shape through our process demonstrates a significant transformation in shock wave structure. The shock wave on the upper part of the airfoil has weakened and shifted its position, while the shock wave on the lower part has dissipated. The results of optimization via reinforcement learning clearly indicate that our optimization objectives have been effectively met throughout the iterative process. The question of whether we can leverage the properties of pressure distribution to enhance the efficiency of the optimization process, such as by introducing additional rewards, represents a promising direction for further research. This approach could potentially streamline the optimization process, making it more efficient by directly influencing the aerodynamic characteristics targeted for improvement.
@@
Figure 12: Left: the record curve of lift-drag ratio during the optimization process. Right: the mesh around the optimized airfoil.
6.2 Maximize lift-drag ratio optimization
Figure 13: Left: the mesh generated from DWR-based adaptation. Right: the Mach isoline of the optimized airfoil.
We conducted the optimization task of maximizing the lift-drag ratio of NACA0012 with Mach number $0.8$ and attack angle $1.25^{\circ}$ as well. The value of the lift-drag ratio gets raised from $16.488$ to $48.916$. Optimizing for the lift-drag ratio poses greater challenges than solely focusing on drag due to the complex calculation of lift. Firstly, the calculation of lift is more complicated compared with drag. Evidence from our previous experiments and other researchers [10] have shown that the convergence curve of lift during the calculation is oscillatory. This may be caused by the weaker regularity of the dual solution for the lift coefficient. Even though dual consistency has been considered, such a problem has not been fully figured out. Secondly, the DWR-based mesh adaptation method can not be generalized to the lift-drag ratio easily. Then, the accuracy of the target functional gets influenced a lot. Multi-mesh method[52] can be applied to solve the lift and drag respectively. However, it will bring additional difficulties to the PDE solver.
Currently, we pay more attention to the drag calculation since the denominator demands higher accuracy in a fractional calculation. The mesh generated from DWR-based adaptation as shown in Figure 14 is adopted to calculate the lift-drag ratio. Our experiments, illustrated in Figure 13, highlight potential issues with data anomalies that can arise when the framework for solving the designated functional is not sufficiently accurate. Such inaccuracies can lead to erroneous signals that adversely affect the policy update of the agent. Moreover, it is evident that there is a need for a more task-specific design of the reward functional. The optimization process depicted in Figure 13 underwent an extensive period of exploration to identify a viable direction for improvement. Although the optimized shape has a lift-drag ratio nearly three times than the original one, it seems this airfoil shape does not represent the intended airfoil design. The anticipated shape could potentially exhibit even greater deviations. Thus, the agent should explore more potential shapes to generate the best geometry, which will consume time significantly longer than the current experiment. While the current PDE solver has already been a highly efficient one, an increase of magnitude with respect to the wall time is tough. Multi-agent exploration framework should be constructed in the future to enhance this algorithm. Shock waves present both on the lower and upper side of the airfoil from the result in Figure 14 which may be attributed to the problem that the value of the lift-drag ratio during the optimization process stopped around $45$. In order to develop a task-specific design algorithm for the lift-drag calculation, we need to pay attention to the pressure distribution along the geometry as well. Besides, the analysis should be developed further to ensure the equivalence of the optimization objective.
Refer to captionRefer to caption
Figure 14: Left: the Mach isoline of the initial airfoil. Right: the Mach isoline of the optimized airfoil.

### v2

@@ 4.3 Reward function
In the context of shape optimal design, a well-designed reward function can guide the reinforcement learning algorithm toward more effective and meaningful solutions, ensuring the learned policies are not only efficient but also aligned with the intended outcomes. Even though techniques introduced during the reinforcement learning process, such as discount factors, might prevent strict equivalence in real computational scenarios, maintaining this equivalence as closely as possible facilitates the learning process. Besides, the mechanism is of vital importance to guarantee a precise value of reward where the agent can learn effective policy.
In [27, 15, 50], different reward functions are adopted to meet the requirements of different optimization tasks. It should be noted that the reward function should be carefully designed to support the learning process. A misalignment between the learning objectives and the actual goals, while technically successful within the learning framework, can lead to solutions that cannot fulfill the intended purpose. The resulting policies might only be effective in specific environment and perform poorly when faced with new scenarios. To overcome above issues, we design reward function and analyze the equivalence of the optimization result.
Let us consider the objective of generating a shape with the global minimum drag value. Denote drag as $D(\{sp^{t}_{i}\}_{i=0}^{L+1})$, where $\{sp^{t}_{i}\}_{i=0}^{L+1}$ represents the sample points of airfoil at time step $t$ and $L+2$ is the number of sample points as previously defined. $D(\{sp^{*}_{i}\}_{i=0}^{L+1})$ represents one of the global minimum drag values of an airfoil with corresponding parameters. During the optimization process, a sample points sequence will be obtained, $\{sp^{0}_{i}\}_{i=0}^{L+1},\{sp^{1}_{i}\}_{i=0}^{L+1},\cdots\{sp^{t}_{i}\}_{i=0}^{L+1},\cdots$. While seeking for the objective shape of airfoil, the sequence is expected to converge to a sample points vector equivalently.
Theorem 1 (Equivalence of reward function)
Suppose that $D(\{sp^{*}_{i}\}_{i=0}^{L+1})$ represents the global minimum drag value. If the reinforcement learning process aims to maximize the cumulative reward $Q_{t}=\sum_{s=t}^{+\infty}R_{s}$ for each step $t$ and generate a sample points sequence with the limit $\{sp^{**}_{i}\}_{i=0}^{L+1}$, then the result is one of the global minimum of the optimization objective, i.e., $D(\{sp^{*}_{i}\}_{i=0}^{L+1})=D(\{sp^{**}_{i}\}_{i=0}^{L+1})$.
Proof.   As $D(\{sp^{*}_{i}\}_{i=0}^{L+1})$ represents the global minimum drag value of airfoil, then
$\begin{array}[]{cc}&D(\{sp^{*}_{i}\}_{i=0}^{L+1})\leq D(\{sp^{t}_{i}\}_{i=0}^{L+1})\qquad\forall sp^{t}_{i}\in\mathbb{R}^{2},\forall t\in\mathbb{N}\\ &\{sp^{*}_{i}\}_{i=0}^{L+1}\equiv\{sp^{t_{0}}_{i}\}_{i=0}^{L+1}\quad\text{if\leavevmode\nobreak\ }D(\{sp^{t_{0}}_{i}\}_{i=0}^{L+1})\leq D(\{sp^{*}_{i}\}_{i=0}^{L+1}).\end{array}$ (20)
Since $\{sp^{**}_{i}\}_{i=0}^{L+1}$ denotes the sample points of the airfoil that yield the maximal reward of the learning process, then
$D_{0}+(-D(\{sp^{**}_{i}\}_{i=0}^{L+1}))=\lim\limits_{t\to+\infty}\sum_{s=1}^{t}R_{s}\geq D_{0}+(-D(\{sp^{t}_{i}\}_{i=0}^{L+1}))\qquad\forall sp^{t}_{i}\in\mathbb{R}^{2},\forall t\in\mathbb{N}.$ (21)
@@
The process above demonstrates the equivalence of the reinforcement optimization objective to the primary task. A similar analysis, such as maximizing the lift-drag ratio, can also be derived in this manner.
@@
To further refine the optimization strategy, regularization terms can be added to the reward function. These terms are designed to encourage exploration. By incorporating exploration-driven regularization, the reward function can be represented as:
$\tilde{R}_{t}=(D_{t-1}-D_{t})+\lambda^{\epsilon}_{t}(D_{0}-D_{t}),$ (24)
@@
where $\lambda^{\epsilon}_{t}$ is a weighting coefficient that balances the original objective with the exploration incentive. When the current shape yields an improvement in the objective function value, the agent gets a reward. Conversely, a penalty is incurred for generating a geometry that does not meet the expected criteria. The parameter $\lambda^{\epsilon}_{t}$ shall be decreased with the update of noise. In order to guarantee the equivalence above, the coefficient $\lambda^{\epsilon}_{t}$ is set as an exponential decline function.
Theorem 2 (Equivalence of generalized reward function)
If the learning aims to maximize the cumulative reward $\tilde{Q}_{t}=\sum_{s=t}^{+\infty}\tilde{R}_{s}$ and generates a sample points sequence with limit $\{sp^{***}_{i}\}_{i=0}^{L+1}$, the result is one of the global minimum of the optimization objective, i.e., $D(\{sp^{*}_{i}\}_{i=0}^{L+1})=D(\{sp^{***}_{i}\}_{i=0}^{L+1})$.
Proof.   Suppose that the learning process generate a sequence $(D^{**}_{t})_{t\in\mathbb{N}}$ which has a limit $D(\{sp^{**}_{i}\}_{i=0}^{L+1})$, if the objective is to maximize the cumulative reward $\sum_{s=t}^{+\infty}{R}_{s}.$ Similarly, a sequence $(D^{***}_{t})_{t\in\mathbb{N}}$ is generated with the objective to maximize the cumulative reward $\sum_{s=t}^{+\infty}\tilde{R}_{s}.$ Firstly, the cumulative reward is finite since the coefficient is an exponential decline function, i.e.,
$\sum_{s=1}^{+\infty}\tilde{R}_{s}=\sum_{s=1}^{+\infty}{R}_{s}+\sum_{t=1}^{+\infty}\lambda^{\epsilon}_{t}(D_{0}^{***}-D_{t}^{***})\leq D_{0}^{***}+(D_{0}^{***})\sum_{t=1}^{+\infty}\lambda^{\epsilon}_{t}<+\infty.$ (25)
If $D(\{sp^{***}_{i}\}_{i=0}^{L+1})\leq D(\{sp^{**}_{i}\}_{i=0}^{L+1}),$ from the definition and theorem 1, the equivalence of reward function held naturally. To the contrary, suppose $D(\{sp^{***}_{i}\}_{i=0}^{L+1})>D(\{sp^{**}_{i}\}_{i=0}^{L+1})$ and denote $\delta=(D(\{sp^{***}_{i}\}_{i=0}^{L+1})-D(\{sp^{**}_{i}\}_{i=0}^{L+1}))/2$. Then, $\exists N_{\delta}$, s.t., $\forall t>N_{\delta}$, $D_{t}^{***}-D_{t}^{**}>\delta/2$. While the sequence $(D^{***}_{t})_{t\in\mathbb{N}}$ and $(D^{**}_{t})_{t\in\mathbb{N}}$ are generated with maximizing $Q_{t}$ and $\tilde{Q}_{t}$, specifically, we have
$+\infty>\tilde{Q}_{N_{\delta}}=\sum_{t=N_{\delta}}^{+\infty}\tilde{R}_{s}=\sum_{t=N_{\delta}}^{+\infty}\left((D_{t-1}^{***}-D_{t}^{***})+\lambda^{\epsilon}_{t}(D_{0}^{***}-D_{t}^{***})\right)$ (26)
$\geq\sum_{t=N_{\delta}}^{+\infty}\left((D_{t-1}^{**}-D_{t}^{**})+\lambda^{\epsilon}_{t}(D_{0}^{**}-D_{t}^{**})\right)\geq\sum_{t=N_{\delta}}^{+\infty}\left((D_{t-1}^{***}-D_{t}^{***})+\lambda^{\epsilon}_{t}(D_{0}^{**}-D_{t}^{**})\right).$
Since the sequences both starts from $D_{0}$ and share the same coefficient $\lambda^{\epsilon}_{t}$, the following inequality can be derived
$\sum_{t=N_{\delta}}^{+\infty}\lambda^{\epsilon}_{t}(D_{t}^{**}-D_{t}^{***})\geq 0.$ (27)
However, the inequality cannot held since $D_{t}^{***}-D_{t}^{**}>\delta/2,\leavevmode\nobreak\ \forall\leavevmode\nobreak\ t>N_{\delta}$, which implies the equivalence of this generalized reward function is held. $\Box$
Thus this generalized reward function will encourage exploration at the initial period while the result can converge to the optimization objective with the iteration getting processed. These mechanisms ensure a more robust and effective search for the optimal airfoil shape, facilitating the discovery of innovative designs that minimize drag while navigating the complex landscape of possible configurations.
4.4 Twin delayed DDPG algorithm
@@
With the introduced modules, we are going to develop a more robust framework, TD3. It introduces several key improvements that enhance the stability and performance of learning tasks. The primary advantage of TD3 is its twin-critic network architecture. It involves two separate critic networks to estimate the Q-values, where the smaller one is used for the policy update. This approach effectively reduces the overestimation bias often seen in DDPG, leading to more stable and reliable learning. Additionally, TD3 introduces policy update delays, where the policy network is updated less frequently than the critic. This delay helps mitigate the risk of the policy network exploiting inaccuracies in the Q-value estimates, further enhancing the stability of the learning process. With such a more stable framework, we constructed Algorithm 2.
@@
0:  Initial airfoil shape data, actor parameters $\theta^{a}$, critic parameters $\theta^{c_{1}}$ and twin critic parameters $\theta^{c_{2}}$, target networks $\theta^{a}_{target}$, $\theta_{target}^{c_{1}}$ and $\theta_{target}^{c_{1}}$, ReplayBuffer
0:  Trained Twin Delayed DDPG agent
1:  Set the target parameters $\theta^{a}_{target}\leftarrow\theta^{a},\theta^{c_{1}}_{target}\leftarrow\theta^{c_{1}},\theta^{c_{2}}_{target}\leftarrow\theta^{c_{2}}$
2:  for each episode in max_episodes do
3:     if Exploration or reach the minimum threshold then
4:        Reset the environment and get the initial state
5:     else
6:        for step in episode do
7:           if Exploration then
8:              The agent produces a random action for execution
9:           else
10:              The agent selects an action based on $\theta^{a}_{target}$
11:              Noises are added to the action and corresponding parameters
12:              The agent executes the action based on the current state
13:              Update the cache state
14:           end if
15:           Observe the new state and calculate the reward based on AFVM4CFD
16:           if reward $<$ min_reward then
17:              Reload the cache state
18:              Execute the opposite action with corresponding parameters
19:           end if
20:           Save the tuple (state, action, reward, next_state) to the ReplayBuffer
21:           if ReplayBuffer has enough experiences then
22:              Sample experiences (Recent, Best, Random) from ReplayBuffer
23:              Compute the target actions based on $\theta^{a}_{target}$
24:              Compute the target Q-value $\min(Q_{\theta_{target}^{c_{1}}},Q_{\theta_{target}^{c_{2}}})$
25:              Update the critic networks $\theta^{c_{1}}$, $\theta^{c_{2}}$
26:              Soft update critic networks $\theta_{target}^{c_{i}}\leftarrow\tau\theta_{target}^{c_{i}}+(1-\tau)\theta^{c_{i}}$ for $i=1,2$
27:              if $step\equiv 0\pmod{delay\_step}$ then
28:                 Update the actor network $\theta^{a}$
29:                 Soft update actor networks $\theta_{target}^{a}\leftarrow\tau\theta_{target}^{a}+(1-\tau)\theta^{a}$
30:              end if
31:           end if
32:           Update the state
33:        end for
34:     end if
35:  end for
@@ 5 Numerical Result
5.1 Drag minimization
The initial example focus on minimizing drag calculations. The study utilizes a NACA0012 airfoil, positioned within a circular domain with a radius of $35$, under a boundary condition of mirror reflection. The operational conditions are set with Mach number $0.85$ and attack angle $0^{\circ}$.
Figure 9: Record curves during the optimization. Left: noise coefficient. Right: drag value.
To ensure the integrity of the airfoil’s structure during training, a constraint is imposed, mandating a minimum thickness of 0.1 for sections where $x<0.1$ and $x>0.7$. Should a potential action result in an airfoil design adhering to this thickness constraint, it is deemed physically infeasible. Consequently, this action will not be executed, leading to a negative reward.
As shown in Algorithm 3, the agent embarks on accumulating experience through random actions performed four times, each starting from the initial configuration across $64$ time steps. The initial recorded drag value is $0.0452106$.
Refer to captionRefer to captionRefer to caption
Figure 10: Right: initial airfoil, NACA0012. Middle: the airfoil generated from the 15th episode. Right: the result of the optimized airfoil.
Training starts with a moderate batch size of $256$, spanning a total of $75$ epochs, with each epoch comprising 8 time steps. The noise coefficient gets updated every three epochs, following an exponential decline. As illustrated in Figure 9, the drag value exhibits a decrement trend as iterations advance. Even though the record curve is not smooth as shown in the record line, it still evolved towards the optimization objective. The non-smooth phenomenon can be explained by the fact of noise. Even if the potential action from the actor neural network generates an effective action with corresponding parameters, the noise may distract the agent from the potential actions. However, such a noise module is indispensable to help the agent explore the possibility of different states.
Throughout the optimization process, the agent explores various airfoil shapes. For instance, Figure 10 exhibits a configuration from the $15^{th}$ epoch, with a drag value of $0.031537$. The drag value’s trajectory halts around $0.02$, constrained by the minimum thickness. Encounters with this restriction trigger a negative reward, arousing exploration toward alternate states. Notably, as the exploration progresses, particularly in the concluding epoch, the airfoil’s tail approaches the thickness restriction.
The interconnection between drag optimization and shock waves is a fundamental aspect of airfoil design, where minimizing the impact of shock waves is pivotal for reducing wave drag. This concept has been supported and observed across various studies within the field [51]. From our experiment, the NACA 0012 airfoil initially exhibited strong shock waves on both the upper and lower surfaces, typically associated with substantial pressure discontinuities near these regions, significantly contributing to increased drag. Through the optimization process, notable changes in the structure of these shock waves were observed. Specifically, the optimization resulted in the attenuation and positional shift of the shock wave on the airfoil’s upper surface, along with the complete dissipation of the lower surface’s shock wave.
@@ 5.1 Drag minimization
Figure 11: Mach isoline. Left: the initial airfoil. Right: optimized airfoil.
5.2 Lift-drag ratio maximization
Figure 12: Left: record curve of lift-drag ratio during the optimization process. Right: mesh around the optimized airfoil.
We conducted the optimization task of maximizing the lift-drag ratio of NACA0012 with Mach number $0.8$ and attack angle $1.25^{\circ}$. The value of the lift-drag ratio gets raised from $16.488$ to $48.916$. Optimizing for the lift-drag ratio poses greater challenges than solely focusing on drag due to the singularity of the objective functional. Firstly, the calculation of lift is more complicated than drag. Evidence from our previous experiments and other researchers [8] have shown that the convergence curve of lift during the calculation is oscillatory. This may be caused by the weaker regularity of the dual solution for the lift coefficient. Secondly, the DWR-based mesh adaptation method can not be generalized to the lift-drag ratio easily. Then, the accuracy of the target functional gets influenced a lot. Multi-mesh method[52] can be applied to solve the lift and drag respectively. However, it will bring additional difficulties to the PDE solver.
@@ 5.2 Lift-drag ratio maximization
Figure 13: Left: mesh generated from DWR-based adaptation. Right: Mach isoline.
Currently, we pay more attention to the drag calculation since the denominator demands higher accuracy in the lift-drag ratio calculation. The mesh generated from DWR-based adaptation as shown in Figure 13 is adopted to calculate the lift-drag ratio. Our experiments, illustrated in Figure 12, highlight data anomalies issue that can arise when the framework for solving the objective functional is not sufficiently accurate. Such inaccuracies can lead to erroneous signals that affect the policy update process. It indicates that a more task-specific design of the reward functional need to be designed. The optimization process, depicted in Figure 12, shows an extensive period of exploration to identify a viable direction for improvement. Although the optimized shape has a lift-drag ratio nearly three times than the original one, the airfoil shape does not represent the intended airfoil design. The agent should explore more potential shapes to generate the expected geometry, which will consume time much longer than the current experiment. While the current PDE solver has already been a highly efficient one, an increase of magnitude with respect to the wall time is tough. Multi-agent exploration framework should be constructed in the future to enhance this algorithm. Shock waves can be found in Figure 13, which may explain the reason that the lift-drag ratio stopped around $45$ during the optimization process. To develop a task-specific design algorithm for the lift-drag calculation, we need to pay attention to the pressure distribution along the geometry as well.